OpenStudy (anonymous):

Please help ): Determine the type of boundary line and shading for the graph of the inequality -3x - y > -6 Dashed line with shading on the side that includes the origin. Solid line with shading on the side that does not include the origin. Dashed line with shading on the side that does not include the origin. Solid line with shading on the side that includes the origin.

OpenStudy (anonymous):

i dont get how to do this at all help?

OpenStudy (mathstudent55):

First, change the inequality into its corresponding equation by simply using the = sign instead of the > sign.

OpenStudy (mathstudent55):

Then, graph the line. Since the original problem has a simple > sign, not a >= sign, the line will be dashed.

ganeshie8 (ganeshie8):

another trick to figure out the origin thing is to : 1. plugin x=0, y=0. if the inequality satisfies, then it includes origin.

ganeshie8 (ganeshie8):

-3x - y > -6 plugin x = 0, y = 0 -3(0) - 9 > -6 0 > -6 true

ganeshie8 (ganeshie8):

So, it includes origin

OpenStudy (anonymous):

i still dont get it but thank you @ganeshie8

ganeshie8 (ganeshie8):

oops sorry !! i made a mistake in my calculation above. that wud have confused u

ganeshie8 (ganeshie8):

it must be like this :- -3x - y > -6 plugin x = 0, y = 0 -3(0) - 0 > -6 0 > -6 true

ganeshie8 (ganeshie8):

its okay if u dont get it now, do two more problems like these... im sure these wud make sense to you :) good luck !

OpenStudy (anonymous):

so the answer is A then correct?

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

A is correct !!!

ganeshie8 (ganeshie8):

dashed line cuz there is no = sign in the given inequality

ganeshie8 (ganeshie8):

includes origin, cuz it satisfies when we plugged x=0, y=0

OpenStudy (anonymous):

will you help me with another please? @ganeshie8

ganeshie8 (ganeshie8):

sure post it :)

OpenStudy (anonymous):

thank you! and okay